OpenStudy (anonymous):

Factor completely p4 – 16

OpenStudy (anonymous):

(p2 + 4)(p + 4)(p – 4) (p2 + 2)(p2 + 8) (p2 – 4)2 (p2 + 4)(p + 2)(p – 2)

OpenStudy (anonymous):

options^

OpenStudy (anonymous):

((p^2)+4)((p^2)-4)= Simplify this

hartnn (hartnn):

do u know the formula a^2-b^2=(a-b)(a+b)??

OpenStudy (anonymous):

Use the difference of squares formula x2−a2=(x−a)(x+a) and note that x4=(x2)2 if you do this correctly you should be able to do it twice.

OpenStudy (ghazi):

p^4= (p^2)^2 ...so you'll have (P^2-4)(P^2+4)...that is (a^2-b^2)= (a+b)(a-b)

OpenStudy (anonymous):

I originally got (p^2+4) (p^2-4) but that's not one of the options :(

OpenStudy (anonymous):

(p-2)(p+2)(p^2+4)

OpenStudy (ghazi):

further you can break P^2-4= (p+4)(p-4)

OpenStudy (anonymous):

Brittany, your answe cna be simplified as I said before.

OpenStudy (anonymous):

That's it Ghazi! *Takes out shotgun* Get of my lawn you hooligan.

OpenStudy (anonymous):

(p2 + 4)(p + 2)(p – 2)

OpenStudy (anonymous):

Of course, I jest.

OpenStudy (anonymous):

so it's the last choice then? I simplified it

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

yes like i said

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

welcome